Sol Summary – May 16th

Crew 281 Sol Summary Report 16-05-2023
Sol: 2
Summary Title: Around the Hab and Experiments
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
Before the EVA prep began Megan Kane (Commander) checked the greenhab.
EVA 3: Ritupriya Patil (XO, EVA Lead), Ana Pires (Crew Scientist)(EVA Comm), Rachel Jones (HSO). Hab Comm: Megan Kane
EVA prep began at 7:45 am when we started prepping our suits, radio comms, helmets and air circulation. KC Shasteen (Crew Engineer) made the pre-EVA checks with radios and spacesuit. We donned the EVA team’s suits, conducted a 15 minute pre-breath followed by a 5 minute depressurization in the air lock, leaving the airlock at 8:35 am. Waterchecks were scheduled for every 45 minutes. Mid EVA one suit required adjustments. All objectives were completed. Due to the suit requiring further adjustment the EVA crew returned to the hab early at 0956. After the EVA crew returned, KC and Megan assisted with adjusting the suit for future EVAs and, hopefully, solving the issue with the fit. Additionally during the EVA there was a communications issue that was successfully resolved.
In the Greenhab an airflow diverter was added to the fan to improve the conditions for the plants.
In the science dome, 2 research projects were worked on.

Look Ahead Plan:
Tomorrow we plan to conduct the first Geotechnical survey for the Crew Scientists research.
Additional experiments are in the works and existing experiments will be continued.
Anomalies in work: None
Weather: Fine
Crew Physical Status:
· 1 crew in suboptimal condition, monitoring
EVA: Completed 1 today
Reports to be filed:
● Greenhab Report
● EVA Reports EVA 3
● Ops Report
● Commander’s Report
Support Requested:
● Issue with power system was observed and reported to Mission Support.
● Trash pickup requested for this evening.

Copyright © The Mars Society. All rights reserved. | Main Site